A new project has an initial cost of $148,000. The equipment will be depreciated on a straight-line basis to a book value of $49,000 at the end of the four-year life of the project. The projected net income each year is $14,500, $17,700, $22,600, and $14,400, respectively. What is the average accounting return?

  • A. 17.56%

  • b.21.10%

  • c.18.82%

  • d.23.38%

  • E.9.35%
Answer #1

Answer: A.17.56%

Explanation

Average accounting return = Average annual net income/ Average investment

Average investment = (book value at year 1+ book value at end of useful life) / 2

= (148,000+49,000)/2 = $98,500

Average annual net income = total net income over investment period/number of years

=(14,500+17,700+22,600+14,400)/4 = $17,300

Average accounting return = $17,300/$98,500 = 0.1756 or 17.56%